Four unit circles pass through the origin and have their centres on the coordinate axis.What is the area of the quadrilateral whose vertices are the point of intersection(in pairs) of the circles?

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago
Dear aniket equation of circles are (x-1)^2 + y^2 =1 (x+1)^2 + y^2 =1 (y-1)^2 + x^2 =1 (y+1)^2 + x^2 =1 point of intersection are (1,1) , (1,-1) , (-1,1) and (-1,-1) this the vertex of square of side 2 so area = 2*2 =4 Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE preparation. All the best. Regards, Askiitians Experts Badiuddin